# Focal Surface Holographic Light Transport using Learned Spatially Adaptive Convolutions

### v1

@@ 2.1. Spatially Adaptive Convolution
where $\tilde{c}$ and $\hat{c}$ indicate the number of input and output channels. The indices satisfy $1\leq c^{\prime}\leq\tilde{c}$ and $1\leq c\leq\hat{c}$. The pair $(x^{\prime},y^{\prime})$ belongs to the set $\Omega(k)$, which specifies a $k\times k$ convolutional window. The summation operation acts on all input channels, which implies that each input channel contributes to every output channel. According to Eq. (1), this operation is characterized by a kernel that is spatially shared and content-independent. Learning-based light transport models could use Eq. (1) as a basic operation. However, it is challenging for this method to project a hologram onto a focal surface. As each pixel on the hologram plane may correspond to a different depth on the focal surface, which makes the  kernel a sub-optimal choice to capture spatially varying features (Su et al., 2019), including focusing or out-of-focus effects due to depth variance. A typical solution is to employ a large number of parameters for feature encoding, resulting in an increased memory footprint. Alternatively, we could consider using  convolution (Zheng et al., 2021; Su et al., 2019). The  kernel $\mathbf{V}\in\mathbb{R}^{\hat{c}\times h\times w\times\tilde{c}\times k\times k}$ incorporates two new dimensions $h,w$ into  kernel, where $h$ and $w$ indicate height, and width of the output feature. However, relying solely on  kernels may increase model capacity due to extra parameters, particularly when $h$ and $w$ are large. These alternative designs all demand extra network capacity.
@@
To address these problems, we propose the . Our method reduces the network parameters by multiplying the  kernel with the standard  kernel. Initially, the  kernel $\mathbf{V}\in\mathbb{R}^{1\times h\times w\times\tilde{c}\times k\times k}$ is introduced, the output channel is set to $1$ to reduce the number of parameters. The  kernel $\mathbf{A}\in\mathbb{R}^{\hat{c}\times h\times w\times\tilde{c}\times k\times k}$ is computed by multiplying the $\mathbf{W}$ and $\mathbf{V}$, which defined as:

### v2

@@ 2.1. Spatially Adaptive Convolution
where $\tilde{c}$ and $\hat{c}$ indicate the number of input and output channels. The indices satisfy $1\leq c^{\prime}\leq\tilde{c}$ and $1\leq c\leq\hat{c}$. The pair $(x^{\prime},y^{\prime})$ belongs to the set $\Omega(k)$, which specifies a $k\times k$ convolutional window. The summation operation acts on all input channels, which implies that each input channel contributes to every output channel. According to Eq. (1), this operation is characterized by a kernel that is spatially shared and content-independent. Learning-based light transport models could use Eq. (1) as a basic operation. However, it is challenging for this method to project a hologram onto a focal surface. As each pixel on the hologram plane may correspond to a different depth on the focal surface, which makes the  kernel a sub-optimal choice to capture  features (Zheng et al., 2021; Xu et al., 2020), including focusing or out-of-focus effects due to depth variance. A typical solution is to employ a large number of parameters for feature encoding, resulting in an increased memory footprint. Alternatively, we could consider using  convolution (Zheng et al., 2021; Xu et al., 2020). The  kernel $\mathbf{V}\in\mathbb{R}^{\hat{c}\times h\times w\times\tilde{c}\times k\times k}$ incorporates two new dimensions $h,w$ into  kernel, where $h$ and $w$ indicate height, and width of the output feature. However, relying solely on  kernels may increase model capacity due to extra parameters, particularly when $h$ and $w$ are large. These alternative designs all demand extra network capacity.
@@
To address these problems, we utilize the  based on (Xu et al., 2020). Our method reduces the network parameters by multiplying the  kernel with the standard  kernel. Initially, the  kernel $\mathbf{V}\in\mathbb{R}^{1\times h\times w\times\tilde{c}\times k\times k}$ is introduced, the output channel is set to $1$ to reduce the number of parameters. The  kernel $\mathbf{A}\in\mathbb{R}^{\hat{c}\times h\times w\times\tilde{c}\times k\times k}$ is computed by multiplying the $\mathbf{W}$ and $\mathbf{V}$, which defined as: